# BiVLC: Extending Vision-Language Compositionality Evaluation with Text-to-Image Retrieval

### v1

@@ Abstract
Existing Vision-Language Compositionality (VLC) benchmarks like SugarCrepe are formulated as image-to-text retrieval problems, where, given an image, the models need to select between the correct textual description and a synthetic hard negative text. In this work, we present the Bidirectional Vision-Language Compositionality (BiVLC) dataset. The novelty of BiVLC is to add a synthetic hard negative image generated from the synthetic text, resulting in two image-to-text retrieval examples (one for each image) and, more importantly, two text-to-image retrieval examples (one for each text). Human annotators filter out ill-formed examples ensuring the validity of the benchmark. The experiments on BiVLC uncover a weakness of current multimodal models, as they perform poorly in the text-to-image direction. In fact, when considering both retrieval directions, the conclusions obtained in previous works change significantly. In addition to the benchmark, we show that a contrastive model trained using synthetic images and texts improves the state of the art in SugarCrepe and in BiVLC for both retrieval directions. The gap to human performance in BiVLC confirms that Vision-Language Compositionality is still a challenging problem. BiVLC and code are available at https://imirandam.github.io/BiVLC_project_page.
@@ 1 Introduction
In this paper, we propose a semi-automatic method to generate an evaluation dataset for VLC that includes image-to-text and text-to-image retrieval (referred as I2T and T2I respectively). We refer to this problem bidirectional Vision-Language Compositionality, since it combines both retrieval directions. For that purpose, we extend the SugarCrepe dataset [Hsieh et al., 2024], the reference benchmark for image-to-text retrieval-based VLC. We leverage the hard negative textual descriptions provided in the dataset and, for every hard negative text, we generate four candidate images using off-the-shelf text-to-image generators. We ask human annotators to select the image that better describes the text and to discard those instances where no image is correct. In a second annotation step, we ask annotators to remove instances where the generated image matches both captions, resulting in an ambiguous instance. As a result, we build a high-quality evaluation dataset called BiVLC (Bidirectional Vision-Language Compositionality), with almost three thousand instances composed by two images and two captions, which accounts to more than eleven thousand retrieval instances (Figure 1). BiVLC allows us to evaluate models on both text-to-image and image-to-text retrieval, offering a more complete picture of VLC capabilities. Our analysis will show that distinguishing between synthetic and natural images-texts is not enough to perform well in BiVLC.
By means of extensive experimentation and analyses, we present the following contributions: (i) Humans perform comparably for both retrieval directions, but multimodal models are significantly worse for text-to-image retrieval; (ii) Bidirectional VLC is more difficult than image-to-text retrieval; (iii) The strongest models for image-to-text retrieval are not necessarily the strongest for bidirectional VLC; (iv) Training with hard negative images can boost the performance of multimodal models, obtaining the best results for BiVLC, but still far from humans.
@@ 2 Related Work
To measure Vision-Language Compositionality, image-to-text retrieval is the dominant paradigm in the literature. Different datasets and benchmarks have been proposed following this approach: for example, [Ma et al., 2023] introduced CREPE, a large dataset containing hard negative texts generated by means of heuristic rules. The dataset was designed to measure sistematicity and productivity, two important aspects of compositionality. A similar benchmark is ARO [Yuksekgonul et al., 2022], which stands for Attribution, Relation, and Order, and also uses heuristic rules to produce hard negative captions given a matching image and text. ARO also stands out for its size, with more than 50,000 test cases.
@@
Due to the use of heuristic rules for hard negative text generation, [Hsieh et al., 2024] found that very high accuracies can be obtained for those two datasets, even without using the images at all. Taking advantage of the biases introduced by the rule-based hard negative texts, such as nonsensical phrases or grammatically incorrect texts, purely textual models outperform the best multimodal models. In consequence, [Hsieh et al., 2024] proposed a new methodology to produce hard negative texts, leveraging modern Large Language Models (LLM), adversarial refinement and human annotation. As a result, the SugarCrepe dataset was proposed, which showed that some of the conclusions obtained when using previous datasets, do not hold anymore. In this work, we extend SugarCrepe with hard negative images, i.e. with images that match the hard negative textual captions and differ from the original caption and image in compositional details. This way, we can measure bidirectional composionality, i.e. we can measure the performance of multimodal models for image-to-text retrieval and also for text-to-image retrieval.
@@
To the best of our knowledge, Winoground [Thrush et al., 2022] is the only dataset that also includes both retrieval directions to measure model performance. However, as shown by [Diwan et al., 2022], Winoground has several problems: (i) it is very small (it contains 400 instances), which makes the comparison among models difficult, (ii) it contains very few instances which actually measure Vision-Language Compositionality (only 171), and (iii) some other challenges like commonsense reasoning or locating small, out-of-focus objects in low resolution images are very important to perform well on the task. In this work, mimicking Winoground, we also build every instance of the dataset with two images and two captions, but we only focus on compositionality and we provide almost 3 thousand instances, increasing the size of Winoground over more than 7 times.
@@ 3 BiVLC: Bidirectional Vision-Language Compositionality Dataset
We introduce BiVLC, a new benchmark to measure bidirectional vision-language compositionality. Each instance of our dataset is comprised by two images and two captions, with the first pair labeled as positive (image+ and caption+) and the second pair labeled as negative (image- and caption-). For each of the images, one of the captions is the correct one and the other is a hard negative distractor. The same happens for each of the captions. Thus, we can measure image-to-text and text-to-image retrieval with hard negative pairs. Some examples are depicted in Figure 2 (See Appendix H for more examples).
@@ 3.1 Dataset construction methodology
To generate BiVLC, we take SugarCrepe as the basis, since it already contains around 7 thousand images with their corresponding positive captions and hard negative captions, carefully created and filtered to avoid any undesirable biases. We apply the following steps:
@@
Since the formatting of the captions provided in COCO (which are used also for SugarCrepe) can vary, starting in lower case, all capitals, missing the final punctuation and so on, we have taken two steps to unify the formatting with the generated captions: 1) process all captions so that they start with capitals and continue in lower case and 2) add the final punctuation mark in case it is missing. This way, we avoid any form-based bias between natural and synthetic captions.
@@
Step 3 - Ask to human annotators to choose the best generated image:
Given a hard negative caption from SugarCrepe and the 4 images generated in Step 2, we hired human annotators to choose the best of the images. SDXL-DPO does not always generate a suitable image for a given textual description, so annotators were allowed to discard all 4 images. In such cases, we also discard the hard negative caption and its associated positive caption and image (more details about the annotation process in Appendix I.1)). After this annotation step, we have 5,382 instances of two images and two captions. In this step, we used 10 annotators divided into pairs to score a total of 500 annotations and obtain inter-tagger agreement. We calculated Cohen’s Kappa score for each pair (Appendix I.3). Since there can be more than one valid images for a caption, we focused on the agreement between selecting any of the 4 images or none. We obtained a mean score of 0.49 between the 5 groups of annotators, which is interpreted as moderate agreement.
@@
During Step 3, annotators only see the hard negative caption and the four generated images. As they do not see the original image and its caption, there can be cases where the annotators marked an image as suitable for a given caption, but that same image could also match the original caption, creating an ambiguous instance. This is specially problematic for the Add category, for which by definition, the generated image also follows the original caption. To minimize those ambiguous instances, we run a filtering step, hiring again human annotators. More concretely, for each instance we give annotators the original caption and both images. We ask them to choose the image that best describes the caption. If the annotator chooses the generated image, we consider that ambiguous and remove the instance from the dataset. If both images are equally suitable, annotators can mark those instances as ambiguous and we also filter them out.
@@ 3.2 Evaluation metrics
The main task defined in BiVLC consists on, given two images and two captions, selecting which image is paired with which caption. To measure the performance on the task, we use the same metrics as Winoground [Thrush et al., 2022], namely, the text, image and group scores. For clarity and coherence, we rename the text score as image-to-text accuracy (I2T for short) and the image score as text-to-image accuracy (T2I). Basically, I2T measures the performance for image-to-text retrieval, thus it is equivalent to SugarCrepe. However, in contrast with SugarCrepe, for each instance in our dataset, we actually have two image-to-text retrieval examples. To obtain a perfect I2T score, the correct captions for both images have to be selected. The T2I is similarly defined for text-to-image retrieval. Finally, the group score is the main metric, since it combines the performance for image-to-text and text-to-image retrieval. To obtain a perfect group score for a given instance, both images have to be matched with the suitable captions and both captions with the suitable images (formal definitions can be found in Appendix C).
@@ 4 Experiments and Findings in BiVLC
In this section, we show the experiments performed in BiVLC and our findings using state-of-the-art open multimodal models for VLC. More concretely, we select the following models: (i) the original CLIP contrastive model [Radford et al., 2021], trained on a private dataset of 400M image-caption pairs; (ii) CLIPCOCO, a CLIP model we fine-tuned on the COCO dataset [Lin et al., 2014] using in-batch negatives (training details in Appendix E); (iii) NegCLIP, another CLIP model fine-tuned on the COCO dataset augmented with rule-based hard negative captions [Yuksekgonul et al., 2022]; (iv) Generative Negative Mining (GNM) [Sahin et al., 2024], a CLIP model fine-tuned on a dataset which contains hard negative texts and images derived from COCO using image editing techniques. For all of them, we use ViT-B/32 as the backbone visual model, to ensure fair comparisons (cf Appendix E).
@@
Table 2: Results for SugarCrepe and BiVLC. For the later, we provide the three metrics I2T, T2I and Group score. Human and random performance are also depicted.
Model SugarCrepe BiVLC
@@
Human 98.93 90.40 93.00 86.80
Random 50.00 25.00 25.00 16.67
CLIP 76.56 75.83 52.40 49.06
@@
While humans show a balanced performance between text-to-image and image-to-text retrieval scores, the models behave very differently, i.e. the T2I score is in all cases more than 20 points lower than the I2T score. Those results show that current multimodal models do not have human comparable VLC skills at all and highlight the importance of measuring text-to-image retrieval.
@@
The gap between the best model and humans in BiVLC is almost 26 points (based on group score, since it is the metric that measures bidirectional performance); for SugarCrepe, 13 points. This means that the bidirectional task is comparatively more difficult than image-to-text retrieval for models than for humans and that there is more room for improvement.
@@
The rank of the multimodal models is different in each of the datasets. More concretely, while NegCLIP is the best model for SugarCrepe, the best model in BiVLC is CLIPCOCO, which uses COCO training examples as in-batch negatives. This is surprising: while NegCLIP uses hard negative texts for training and GNM also adds hard negative images (both derived from COCO), CLIPCOCO does not use any technique to mine any kind of hard negatives. This is very important, since it means that, when considering bidirectional VLC, all the special training approaches designed to enhance VLC so far, are not better than just using the COCO dataset itself.
@@ 5 Exploring training strategies for vision-language compositionality
There are two main strategies in the literature to improve the VLC skills of a multimodal model: (i) using hard negative texts for training (exemplified by NegCLIP [Yuksekgonul et al., 2022]), and (ii) using both, hard negative texts and images (the case of GNM [Sahin et al., 2024]). None of them has been successful in BiVLC. In Sections 5.1 and 5.2 we analyse the limitations of current implementations and propose two new models based on those strategies. We then evaluate them in SugarCrepe and BiVLC (Section 5.3).
@@ 5.1 TROHN-Text: Training on Hard Negative Texts
We used the open-source LLM OpenChat-3.5-0106 [Wang et al., 2023] and the templates provided in SugarCrepe (more details in Appendix F). We generated 591,753 negative captions for each category, for a total of 4,142,271. We then defined a heuristic to filter out the generations that do not follow the templates. After this filtering, we have a total of 3,652,846 negative captions. Finally, we divided the instances consisting of two captions and one image randomly into 80% for training and 20% for validation (more details can be found in Appendix B). We fine-tune a pretrained CLIP model on TROHN-Text and call it CLIPTROHN-Text (training details in Appendix E).

### v2

@@ Abstract
Existing Vision-Language Compositionality (VLC) benchmarks like SugarCrepe are formulated as image-to-text retrieval problems, where, given an image, the models need to select between the correct textual description and a synthetic hard negative text. In this work, we present the Bidirectional Vision-Language Compositionality (BiVLC) dataset. The novelty of BiVLC is to add a synthetic hard negative image generated from the synthetic text, resulting in two image-to-text retrieval examples (one for each image) and, more importantly, two text-to-image retrieval examples (one for each text). Human annotators filter out ill-formed examples ensuring the validity of the benchmark. The experiments on BiVLC uncover a weakness of current multimodal models, as they perform poorly in the text-to-image direction. In fact, when considering both retrieval directions, the conclusions obtained in previous works change significantly. In addition to the benchmark, we show that a contrastive model trained using synthetic images and texts significantly improves over the base model in SugarCrepe and in BiVLC for both retrieval directions. The gap to human performance in BiVLC confirms that Vision-Language Compositionality is still a challenging problem. BiVLC and code are available at https://imirandam.github.io/BiVLC_project_page.
@@ 1 Introduction
In this paper, we propose a semi-automatic method to generate an evaluation dataset for VLC that includes image-to-text and text-to-image retrieval (referred as I2T and T2I respectively). We refer to this problem bidirectional Vision-Language Compositionality, since it combines both retrieval directions. For that purpose, we extend the SugarCrepe dataset [Hsieh et al., 2024], the reference benchmark for image-to-text retrieval-based VLC. We leverage the hard negative textual descriptions provided in the dataset and, for every hard negative text, we generate four candidate images using off-the-shelf text-to-image generators. We ask human annotators to select the image that better describes the text and to discard those instances where no image is correct. In a second annotation step, we ask annotators to remove instances where the generated image matches both captions, resulting in an ambiguous instance. As a result, we build a high-quality evaluation dataset called BiVLC (Bidirectional Vision-Language Compositionality), with almost three thousand instances composed by two images and two captions, which accounts to more than eleven thousand retrieval instances (Figure 1). In addition to the human filtering, we performed experiments which show that distinguishing between synthetic and natural images-texts is not enough to perform well in BiVLC, confirming the validity of the synthetic images.
BiVLC allows us to evaluate models on both text-to-image and image-to-text retrieval, offering a more complete picture of VLC capabilities. The main contribution of this work is the dataset itself, as well as our findings based on the experiments and analysis on BiVLC. We summarize those findings as follows: (i) Humans perform comparably for both retrieval directions, but multimodal models are significantly worse for text-to-image retrieval; (ii) Bidirectional VLC is more difficult than image-to-text retrieval; (iii) The strongest models for image-to-text retrieval are not necessarily the strongest for bidirectional VLC; (iv) Training with hard negative images can boost the performance of multimodal contrastive models, obtaining competitive results for BiVLC, but still far from humans.
@@ 2 Related Work
To measure Vision-Language Compositionality, image-to-text retrieval is the dominant paradigm in the literature. Different datasets and benchmarks have been proposed following this approach: for example, [Ma et al., 2023] introduced CREPE, a large dataset containing hard negative texts generated by means of heuristic rules. The dataset was designed to measure systematicity and productivity, two important aspects of compositionality. A similar benchmark is ARO [Yuksekgonul et al., 2022], which stands for Attribution, Relation, and Order, and also uses heuristic rules to produce hard negative captions given a matching image and text. ARO also stands out for its size, with more than 50,000 test cases.
@@
Due to the use of heuristic rules for hard negative text generation, [Hsieh et al., 2024] found that very high accuracies can be obtained for those two datasets, even without using the images at all. Taking advantage of the biases introduced by the rule-based hard negative texts, such as nonsensical phrases or grammatically incorrect texts, purely textual models outperform the best multimodal models. Similar language biases have also been studied for concurrent benchmarks by [Lin et al., , Tschannen et al., 2024]. In consequence, [Hsieh et al., 2024] proposed a new methodology to produce hard negative texts, leveraging modern Large Language Models (LLM), adversarial refinement and human annotation. As a result, the SugarCrepe dataset was proposed, which showed that some of the conclusions obtained when using previous datasets, do not hold anymore. In this work, we extend SugarCrepe with hard negative images, i.e. with images that match the hard negative textual captions and differ from the original caption and image in compositional details. This way, we can measure bidirectional compositionality, i.e. we can measure the performance of multimodal models for image-to-text retrieval and also for text-to-image retrieval.
@@
In a similar fashion, Winoground [Thrush et al., 2022] also includes both retrieval directions to measure model performance. However, as shown by [Diwan et al., 2022], Winoground has several problems: (i) it is very small (it contains 400 instances), which makes the comparison among models difficult, (ii) it contains very few instances which actually measure Vision-Language Compositionality (only 171), and (iii) some other challenges like commonsense reasoning or locating small, out-of-focus objects in low resolution images are very important to perform well on the task. In this work, mimicking Winoground, we also build every instance of the dataset with two images and two captions, but we only focus on compositionality and we provide almost 3 thousand instances, increasing the size of Winoground considerably.
Recently, two other bidirectional datasets have been published: EqBen [Wang et al., 2023b] and Cola [Ray et al., 2024]. EqBen has been derived from video-text datasets and also offers a set of synthetic images generated with a graphic engine. However, the test set contains low-quality images (e.g. very dark or blurry pictures) [Lin et al., 2024] and authors have released a higher quality subset of 280 pairs of image-text pairs, which is even smaller than Winoground. Cola, on the other hand, has no natural texts, as the captions are produced using templates, and covers only a small subset of compositionality phenomena, as it focuses on object-attributes and spatial relations.
@@ 3 BiVLC: Bidirectional Vision-Language Compositionality Dataset
We introduce BiVLC, a new benchmark to measure bidirectional vision-language compositionality. Each instance of our dataset is comprised by two images and two captions, with the first pair labeled as positive (image+ and caption+) and the second pair labeled as negative (image- and caption-). For each of the images, one of the captions is the correct one and the other is a hard negative distractor. The same happens for each of the captions. Thus, we can measure image-to-text and text-to-image retrieval with hard negative pairs. Some examples are depicted in Figure 2 (See Appendix G for more examples).
@@ 3.1 Dataset construction methodology
To generate BiVLC, we take SugarCrepe as the basis, since it already contains around 7 thousand images with their corresponding positive captions and hard negative captions, carefully created and filtered to avoid any undesirable biases. We apply the following steps (see Figure 3):
Figure 3: Diagram of dataset construction: Starting from SugarCrepe instances, uniformly format positive and hard negative captions (Step 1), generate hard negative images (Step 2), ask human annotators to choose the best generated image (Step 3), and filter out ambiguous instances (Step 4). As a result. we get BiVLC instances, consisting of 2 captions and 2 images.
@@
Since the formatting of the captions provided in COCO (which are used also for SugarCrepe) can vary, starting in lower case, all capitals, missing the final punctuation and so on, we have taken two steps to unify the formatting with the generated captions: 1) process all captions so that they start with capitals and continue in lower case, preserving capitals for named entities, and 2) add the final punctuation mark in case it is missing. This way, we avoid any form-based bias between natural and synthetic captions.
@@
Step 3 - Ask human annotators to choose the best generated image:
Given a hard negative caption from SugarCrepe and the 4 images generated in Step 2, we hired human annotators to choose the best of the images. SDXL-DPO does not always generate a suitable image for a given textual description, so annotators were allowed to discard all 4 images. In such cases, we also discard the hard negative caption and its associated positive caption and image (more details about the annotation process in Appendix H.1)). After this annotation step, we have 5,382 instances of two images and two captions. In this step, we used 10 annotators divided into pairs to score a total of 500 annotations and obtain inter-tagger agreement. We calculated Cohen’s Kappa score for each pair (Appendix H.3). Since there can be more than one valid images for a caption, we focused on the agreement between selecting any of the 4 images or none. We obtained a mean score of 0.49 between the 5 groups of annotators, which is interpreted as moderate agreement.
@@
During Step 3, annotators only see the hard negative caption and the four generated images. As they do not see the original image and its caption, there can be cases where the annotators marked an image as suitable for a given caption, but that same image could also match the original caption, creating an ambiguous instance. This is specially problematic for the Add category, for which by definition, the generated image also follows the original caption. To minimize those ambiguous instances, we run a filtering step, hiring again human annotators. More concretely, for each instance, we give annotators the original caption and both images. We ask them to choose the image that best describes the caption. If the annotator chooses the generated image, we consider that ambiguous and remove the instance from the dataset. If both images are equally suitable, annotators can mark those instances as ambiguous and we also filter them out.
@@ 3.2 Evaluation metrics
The main task defined in BiVLC consists on, given two images and two captions, selecting which image is paired with which caption. To measure the performance on the task, we use the same metrics as Winoground [Thrush et al., 2022], namely, the text, image and group scores. As shown by [Lin et al., ], those evaluation metrics discourage blind solutions and are more robust to language biases. For clarity and coherence, we rename the text score as image-to-text accuracy (I2T for short) and the image score as text-to-image accuracy (T2I). Basically, I2T measures the performance for image-to-text retrieval, thus it is equivalent to SugarCrepe. However, in contrast with SugarCrepe, for each instance in our dataset, we actually have two image-to-text retrieval examples. To obtain a perfect I2T score, the correct captions for both images have to be selected. The T2I is similarly defined for text-to-image retrieval. Finally, the group score is the main metric, since it combines the performance for image-to-text and text-to-image retrieval. To obtain a perfect group score for a given instance, both images have to be matched with the suitable captions and both captions with the suitable images (formal definitions can be found in Appendix D).
@@ 4 Experiments and Findings in BiVLC
In this section, we show the experiments performed in BiVLC and our findings using state-of-the-art open multimodal models for VLC. More concretely, we select the following contrastive models: (i) the original CLIP model [Radford et al., 2021], trained on a private dataset of 400M image-caption pairs; (ii) CLIPCOCO, a CLIP model we fine-tuned on the COCO dataset [Lin et al., 2014] using in-batch negatives (training details in Appendix E); (iii) NegCLIP, another CLIP model fine-tuned on the COCO dataset augmented with rule-based hard negative captions [Yuksekgonul et al., 2022]; (iv) Generative Negative Mining (GNM) [Sahin et al., 2024], a CLIP model fine-tuned on a dataset which contains hard negative texts and images derived from COCO using image editing techniques. For all of them, we use ViT-B/32 as the backbone visual model, to ensure fair comparisons (cf Appendix E). We also evaluate the following generative models: (i) Open CapPa [^1] Code and models in https://wandb.ai/craiyon/cappa-jax/reports/CapPa-Training-vision-models-as-captioners—Vmlldzo4NDUyNDUz, a recent open-source implementation of CapPa [Tschannen et al., 2024] which has shown similar performance as the original model, based on an encoder-decoder architecture with ViT-L-16 with registers and a decoder of 12 layers; (ii) VQAScore [Lin et al., 2024], based on an architecture which combines a CLIP vision encoder (ViT-L-336) with a Flan-T5 LLM [Chung et al., 2024] of varying sizes (XL refers to 3B and XXL to 11B). Table 2 shows the number of parameters of the different models, with VQAscore being significantly larger.
@@
Table 2: Results for existing models (see text for details) on SugarCrepe and BiVLC. For the later, we provide the three metrics I2T, T2I and Group score. Human and random performance are also depicted. Bold for the best of each model family (contrastive and generative).
Model Params SugarCrepe BiVLC
@@
Human N/A 98.93 90.40 93.00 86.80
Random N/A 50.00 25.00 25.00 16.67
Contrastive CLIP 151M 76.56 75.83 52.40 49.06
@@
While humans show a balanced performance between text-to-image and image-to-text retrieval scores, the models behave very differently, i.e. the T2I score is in all cases significantly lower than the I2T score, being Open CapPa the only exception but with very low performance for both metrics. Those results show that current multimodal models do not have human-comparable VLC skills at all, and highlight the importance of measuring text-to-image retrieval.
@@
The gap between the best model and humans in BiVLC is 10 points (based on group score, since it is the metric that measures bidirectional performance); for SugarCrepe, 5 points. This means that the bidirectional task is comparatively more difficult than image-to-text retrieval for models than for humans and that there is more room for improvement.
@@
The rank of the multimodal models is different in each of the datasets. For contrastive models, while NegCLIP is the best model for SugarCrepe, the best model in BiVLC is CLIPCOCO, which uses COCO training examples as in-batch negatives. This is surprising: while NegCLIP uses hard negative texts for training and GNM also adds hard negative images (both derived from COCO), CLIPCOCO does not use any technique to mine any kind of hard negatives. On the other hand, the generative Open CapPa model is clearly ahead of all contrastive models for SugarCrepe, but it is the worst model in BiVLC. This can be explained by the language bias of its LLM decoder, which can be beneficial in SugarCrepe, but is strongly penalized in our bidirectional setting, as already shown in [Lin et al., ].
@@ 5 Exploring training strategies for vision-language compositionality
We will focus on contrastive models for this exploration, since they are the de-facto standard for retrieval problems and are more manageable for training (state-of-the-art generative models are significantly larger). There are two main strategies in the literature to improve the VLC skills of a contrastive model: (i) using hard negative texts for training (exemplified by NegCLIP [Yuksekgonul et al., 2022]), and (ii) using both, hard negative texts and images (the case of GNM [Sahin et al., 2024]). None of them has been successful in BiVLC. In Sections 5.1 and 5.2 we analyse the limitations of current implementations and propose two new models based on those strategies. We then evaluate them in SugarCrepe and BiVLC (Section 5.3).
@@ 5.1 TROHN-Text: Training on Hard Negative Texts
We used the open-source LLM OpenChat-3.5-0106 [Wang et al., 2023a] and the templates provided in SugarCrepe (more details in Appendix F). We generated 591,753 negative captions for each category, for a total of 4,142,271. We then defined a heuristic to filter out the generations that do not follow the templates. After this filtering, we have a total of 3,652,846 negative captions. Finally, we divided the instances consisting of two captions and one image randomly into 80% for training and 20% for validation (more details can be found in Appendix C). We fine-tune a pretrained CLIP model on TROHN-Text and call it CLIPTROHN-Text (training details in Appendix E).